### v1

@@ 3.1 Convergence analysis
Before we prove the iteration complexity of the FF-CR algorithm, we first give the following definition of $\epsilon$-optimal solution with respect to the gradient norm, which provides a performance measure for how close a point $(\hat{x},\hat{y})$ to $(x^{*},y^{*})$ which is an optimal solution of (1) under convex-concave setting.
Definition 3.
A point $(\hat{x},\hat{y})$ is called an $\epsilon$-optimal solution of (1) with respect to the gradient norm if $\|\nabla f(\hat{x},\hat{y})\|\leqslant\epsilon.$
Note that using $\textsc{gap}(x,y;\beta)\leqslant\epsilon$ as the termination criterion requires some prior information about the optimal solution $z^{*}$, which limits the applicability of the algorithm. Therefore, in the FF-CR algorithm proposed in this section, we use $\|\nabla f(x,y)\|\leqslant\epsilon$ as the termination criterion.
Before giving the specific proof details, we first outline the basic idea of iteration complexity analysis. We first prove that once $D_{t}$ in Step 2 of Algorithm 2 satisfies $D_{t}\geqslant\|z_{0}-z^{*}\|$, we can prove $\|\nabla f(x_{t+1},y_{t+1})\|\leqslant\epsilon$ and the algorithm will stop. Otherwise, $D_{t}\leqslant 4\|z_{0}-z^{*}\|$. Then, we estimate the additional number of iterations for such bounded $D_{t}$, and finally prove the iterative complexity of Algorithm 2.
@@
Lemma 4.
Let $\{z_{k}^{t}\}$ and $\{\bar{z}_{k}^{t}\}$ be sequences generated by Algorithm 2. At the $t$-th iteration, for all $k\geqslant 1$, we have
@@
Proof.
By (34) in Algorithm 2, we get
$(z_{k}^{t})^{*}=$ $~{}\arg\min_{x}\max_{y}f(x,y)+\dfrac{\sigma_{k}^{t}}{2}\|x-\bar{x}_{k}^{t}\|^{2}-\dfrac{\sigma_{k}^{t}}{2}\|y-\bar{y}_{k}^{t}\|^{2}$
$=$ $~{}\arg\min_{x}\max_{y}f(x,y)+\dfrac{\sigma_{k-1}^{t}}{2}\|x-\bar{x}_{k-1}^{t}\|^{2}+\dfrac{\sigma_{k}^{t}-\sigma_{k-1}^{t}}{2}\|x-x_{k-1}^{t}\|^{2}$
$~{}-\dfrac{\sigma_{k-1}^{t}}{2}\|y-\bar{y}_{k-1}^{t}\|^{2}-\dfrac{\sigma_{k}^{t}-\sigma_{k-1}^{t}}{2}\|y-y_{k-1}^{t}\|^{2}$
$=$ $~{}\arg\min_{x}\max_{y}f_{k-1}^{t}(x,y)+\frac{\gamma_{k}^{t}\sigma_{k}^{t}}{2}\|x-x_{k-1}^{t}\|^{2}-\frac{\gamma_{k}^{t}\sigma_{k}^{t}}{2}\|y-y_{k-1}^{t}\|^{2}.$ (40)
By the optimal condition of $(z_{k}^{t})^{*}$ in (3.1), we then have
@@
Under the convex-concave setting, $f_{k-1}^{t}(x,y)+\frac{\gamma_{k}^{t}\sigma_{k}^{t}}{2}\|x-x_{k-1}^{t}\|^{2}-\frac{\gamma_{k}^{t}\sigma_{k}^{t}}{2}\|y-y_{k-1}^{t}\|^{2}$ is a $\gamma_{k}^{t}\sigma_{k}^{t}$ strongly-convex-strongly-concave function. Then, by (41), we have
$~{}\gamma_{k}^{t}\sigma_{k}^{t}\|(z_{k-1}^{t})^{*}-(z_{k}^{t})^{*}\|^{2}$
$\leqslant$ $~{}\langle F_{k-1}^{t}\big{(}(z_{k-1}^{t})^{*}\big{)}+\gamma_{k}^{t}\sigma_{k}^{t}\big{(}(z_{k-1}^{t})^{*}-z_{k-1}^{t}\big{)},(z_{k-1}^{t})^{*}-(z_{k}^{t})^{*}\rangle.$ (42)
@@
$~{}\gamma_{k}^{t}\sigma_{k}^{t}\|(z_{k-1}^{t})^{*}-(z_{k}^{t})^{*}\|^{2}$
$\leqslant$ $~{}\gamma_{k}^{t}\sigma_{k}^{t}\langle(z_{k-1}^{t})^{*}-z_{k-1}^{t},(z_{k-1}^{t})^{*}-(z_{k}^{t})^{*}\rangle$
$=$ $~{}\frac{\gamma_{k}^{t}\sigma_{k}^{t}}{2}\big{(}\|(z_{k-1}^{t})^{*}-z_{k-1}^{t}\|^{2}+\|(z_{k-1}^{t})^{*}-(z_{k}^{t})^{*}\|^{2}-\|z_{k-1}^{t}-(z_{k}^{t})^{*}\|^{2}\big{)},$
@@
$\sigma_{k}^{t}=\sum_{i=1}^{k}\alpha_{i}^{t},~{}\sigma_{k}^{t}\bar{z}_{k}^{t}=\sum_{i=1}^{k}\alpha_{i}^{t}z_{i-1}^{t}.$ (45)
@@
$~{}\sigma_{k}^{t}\|\bar{z}_{k}^{t}-(z_{k}^{t})^{*}\|$
$\leqslant$ $~{}\sigma_{k}^{t}\|z_{k-1}^{t}-(z_{k}^{t})^{*}\|+\sigma_{k-1}^{t}\|\bar{z}_{k-1}^{t}-(z_{k-1}^{t})^{*}\|+\sigma_{k-1}^{t}\|(z_{k-1}^{t})^{*}-z_{k-1}^{t}\|$
$\leqslant$ $~{}(\sigma_{k-1}^{t}+\sigma_{k}^{t})\|z_{k-1}^{t}-(z_{k-1}^{t})^{*}\|+\sigma_{k-1}^{t}\|\bar{z}_{k-1}^{t}-(z_{k-1}^{t})^{*}\|.$
Then, (39) holds since $\sigma_{0}^{t}=0$. ∎
The following lemma gives an upper bound of the gradient norm at the point $z_{k}^{t}$ by $\|z_{i}^{t}-(z_{i}^{t})^{*}\|$.
Lemma 5.
Suppose that Assumption 1 holds. Let $\{z_{k}^{t}\}$ and $\{\bar{z}_{k}^{t}\}$ be sequences generated by Algorithm 2. At the $t$-th iteration, for all $k\geqslant 1$, we have
$\|F(z_{k}^{t})\|\leqslant$ $~{}\big{(}\sqrt{\frac{12}{11}}+\frac{72}{11}\big{)}M_{k}^{t}\|z_{k}^{t}-(z_{k}^{t})^{*}\|^{2}+2\sqrt{\frac{12}{11}}M_{k}^{t}\big{(}\sum_{i=1}^{k-1}\|z_{i}^{t}-(z_{i}^{t})^{*}\|\big{)}\|z_{k}^{t}-(z_{k}^{t})^{*}\|$
$~{}+\sqrt{\frac{12}{11}}M_{k}^{t}\|z_{0}^{t}-(z_{0}^{t})^{*}\|\|z_{k}^{t}-(z_{k}^{t})^{*}\|+\big{(}\sqrt{\frac{12}{11}}+1\big{)}\sigma_{k}^{t}\|z_{k}^{t}-(z_{k}^{t})^{*}\|$
$~{}+\sqrt{\frac{12}{11}}\|DF(z_{0}^{t})\|\|z_{k}^{t}-(z_{k}^{t})^{*}\|+\sigma_{1}^{t}\|z_{0}^{t}-(z_{0}^{t})^{*}\|$
$~{}+\sum_{i=2}^{k}(\sigma_{i-1}^{t}+\sigma_{i}^{t})\|z_{i-1}^{t}-(z_{i-1}^{t})^{*}\|.$ (47)
Proof.
By Step 2(d) in Algorithm 2, we have
@@
Under the convex-concave setting, by the optimal condition of $(z_{k}^{t})^{*}$, we further have
$0\leqslant$ $~{}\langle F_{k}^{t}(\tilde{z}),\tilde{z}-(z_{k}^{t})^{*}\rangle$
$=$ $~{}\langle F_{k}^{t}(\tilde{z})-F_{k}^{t}(z_{k}^{t})-DF_{k}(z_{k})(\tilde{z}-z_{k}^{t}),\tilde{z}-(z_{k}^{t})^{*}\rangle$
$~{}+\langle F_{k}^{t}(z_{k}^{t})+DF_{k}^{t}(z_{k}^{t})(\tilde{z}-z_{k}^{t}),\tilde{z}-(z_{k}^{t})^{*}\rangle$
$\leqslant$ $~{}\frac{M_{k}^{t}}{2}\|\tilde{z}-z_{k}^{t}\|^{2}\|\tilde{z}-(z_{k}^{t})^{*}\|-6M_{k}^{t}\|\tilde{z}-z_{k}^{t}\|\langle\tilde{z}-z_{k}^{t},\tilde{z}-(z_{k}^{t})^{*}\rangle,$
@@
$0\leqslant$ $~{}\frac{1}{2}\|\tilde{z}-z_{k}^{t}\|^{2}+\frac{1}{2}\|\tilde{z}-(z_{k}^{t})^{*}\|^{2}-6\big{(}\|\tilde{z}-z_{k}^{t}\|^{2}+\|\tilde{z}-(z_{k}^{t})^{*}\|^{2}-\|z_{k}^{t}-(z_{k}^{t})^{*}\|^{2}\big{)}$
$=$ $~{}-\frac{11}{2}\|\tilde{z}-z_{k}^{t}\|^{2}-\frac{11}{2}\|\tilde{z}-(z_{k}^{t})^{*}\|^{2}+6\|z_{k}^{t}-(z_{k}^{t})^{*}\|^{2},$
@@
$\|F_{k}^{t}(z_{k}^{t})\|=$ $~{}\|DF_{k}^{t}(z_{k}^{t})(\tilde{z}-z_{k}^{t})+6M_{k}^{t}\|\tilde{z}-z_{k}^{t}\|(\tilde{z}-z_{k}^{t})\|$
$\leqslant$ $~{}\|DF_{k}^{t}(z_{k}^{t})\|\|\tilde{z}-z_{k}^{t}\|+6M_{k}^{t}\|\tilde{z}-z_{k}^{t}\|^{2}$
$\leqslant$ $~{}\sqrt{\frac{12}{11}}\|DF_{k}^{t}(z_{k}^{t})\|\|z_{k}^{t}-(z_{k}^{t})^{*}\|+\frac{72M_{k}^{t}}{11}\|z_{k}^{t}-(z_{k}^{t})^{*}\|^{2}.$ (52)
By combining (3.1) with $DF_{k}^{t}(z_{k}^{t})=DF(z_{k}^{t})+\sigma_{k}^{t}I_{m+n}$ and $\frac{\|DF(z_{k}^{t})-DF(z_{0}^{t})\|}{\|z_{k}^{t}-z_{0}^{t}\|}\leqslant M_{k,i}^{t}\leqslant M_{k}^{t}$ in Step 2(d) of Algorithm 2, we have
$\|F_{k}^{t}(z_{k}^{t})\|\leqslant$ $~{}\sqrt{\frac{12}{11}}\|DF(z_{k}^{t})+\sigma_{k}^{t}I_{m+n}-DF(z_{0}^{t})+DF(z_{0}^{t})\|\|z_{k}^{t}-(z_{k}^{t})^{*}\|$
$~{}+\frac{72M_{k}^{t}}{11}\|z_{k}^{t}-(z_{k}^{t})^{*}\|^{2}$
$\leqslant$ $~{}\sqrt{\frac{12}{11}}M_{k}^{t}\|z_{k}^{t}-z_{0}^{t}\|\|z_{k}^{t}-(z_{k}^{t})^{*}\|+\sqrt{\frac{12}{11}}\big{(}\sigma_{k}^{t}+\|DF(z_{0}^{t})\|\big{)}\|z_{k}^{t}-(z_{k}^{t})^{*}\|$
$~{}+\frac{72M_{k}^{t}}{11}\|z_{k}^{t}-(z_{k}^{t})^{*}\|^{2}.$ (53)
@@
$\|F(z_{k}^{t})\|\leqslant$ $~{}\sqrt{\frac{12}{11}}M_{k}^{t}\|z_{k}^{t}-z_{0}^{t}\|\|z_{k}^{t}-(z_{k}^{t})^{*}\|+\sqrt{\frac{12}{11}}(\sigma_{k}^{t}+\|DF(z_{0}^{t})\|)\|z_{k}^{t}-(z_{k}^{t})^{*}\|$
$~{}+\frac{72M_{k}^{t}}{11}\|z_{k}^{t}-(z_{k}^{t})^{*}\|^{2}+\sigma_{k}^{t}\|z_{k}^{t}-(z_{k}^{t})^{*}\|+\sigma_{k}^{t}\|\bar{z}_{k}^{t}-(z_{k}^{t})^{*}\|.$ (54)
@@
$~{}\|z_{k}^{t}-z_{0}^{t}\|$
$=$ $~{}\|z_{k}^{t}-(z_{k}^{t})^{*}+(z_{k}^{t})^{*}-z_{k-1}^{t}+z_{k-1}^{t}-(z_{k-1}^{t})^{*}+\cdots+z_{1}^{t}-(z_{1}^{t})^{*}+(z_{1}^{t})^{*}-z_{0}^{t}\|$
$\leqslant$ $~{}\|z_{k}^{t}-(z_{k}^{t})^{*}\|+\sum_{i=1}^{k-1}2\|z_{i}^{t}-(z_{i}^{t})^{*}\|+\|z_{0}^{t}-(z_{0}^{t})^{*}\|.$ (55)
By plugging (3.1) and (39) into (3.1), the proof is then completed. ∎
@@
Lemma 6.
Suppose that Assumption 1 holds. Let $\{z_{k}^{t}\}$ and $\{\bar{z}_{k}^{t}\}$ be sequences generated by Algorithm 2. Set
$\sigma_{k}^{t}=\frac{\epsilon}{41D_{t}}4^{k},N_{k}^{t}=\left\lceil\big{(}\frac{33\sqrt{3}8^{3-k}H_{k}^{t}D_{t}}{\sigma_{k}^{t}}\big{)}^{\frac{2}{3}}\right\rceil,$
@@
At the $t$-th iteration, if $D_{t}\geqslant\|z_{0}-z^{*}\|$, then we have $\|F(z_{t+1})\|\leqslant\epsilon$, where $z^{*}=(x^{*},y^{*})$ is an optimal solution of (1).
Proof.
By Step 2(c), $(x_{k}^{t},y_{k}^{t},H_{k}^{t})$ is an output of the LF-CR algorithm within $N_{k}^{t}$ iterations and $f_{k}^{t}$ is Hessian Lipschitz continuous. By substituting $f=f_{k}^{t},z=(z_{k}^{t})^{*}$, $z^{*}=(z_{k}^{t})^{*},z_{0}=z_{k-1}^{t},N(\epsilon)-1=N_{k}^{t},H_{N(\epsilon)-1}=H_{k}^{t}$, and $\bar{z}_{N(\epsilon)-1}=z_{k}^{t}$ into (30), we have
$f_{k}^{t}\big{(}x_{k}^{t},(y_{k}^{t})^{*}\big{)}-f_{k}^{t}\big{(}(x_{k}^{t})^{*},y_{k}^{t}\big{)}\leqslant\frac{33\sqrt{3}H_{k}^{t}\|z_{k-1}^{t}-(z_{k}^{t})^{*}\|^{3}}{(N_{k}^{t})^{\frac{3}{2}}}.$ (56)
Under the convex-concave setting, $f_{k}^{t}$ is $\sigma_{k}^{t}$ strongly-convex-strongly-concave.Then, we further have
$\|z_{k}^{t}-(z_{k}^{t})^{*}\|^{2}$ $\leqslant\frac{33\sqrt{3}H_{k}^{t}\|z_{k-1}^{t}-(z_{k}^{t})^{*}\|^{3}}{\sigma_{k}^{t}(N_{k}^{t})^{\frac{3}{2}}}\leqslant\frac{33\sqrt{3}H_{k}^{t}\|z_{k-1}^{t}-(z_{k-1}^{t})^{*}\|^{3}}{\sigma_{k}^{t}(N_{k}^{t})^{\frac{3}{2}}},$ (57)
where the second inequality is by (38). Next, by (57) and induction, we show that for any $k\geqslant 1$, we have
@@
which impies that $\|z_{k+1}^{t}-(z_{k+1}^{t})^{*}\|\leqslant\frac{1}{8}\|z_{k}^{t}-(z_{k}^{t})^{*}\|$. Next, we further bound the norm $\|F(z_{k}^{t})\|$ in (5) by using (58). By (58) and the choice of $\sigma_{k}^{t}$, we have

### v2

@@ 3.1 Convergence analysis
Before proving the iteration complexity of the FF-CR algorithm, we make the following assumption about $f(x,y).$
Assumption 3
The function $f(x,y)$ is globally Hessian Lipschitz continuous.
We then outline the basic idea of iteration complexity analysis. We first prove that once $D_{t}$ in Step 3 of Algorithm 2 satisfies $D_{t}\geqslant\|z_{0}-z^{*}\|$, we can prove $\|\nabla f(x_{t+1},y_{t+1})\|\leqslant\epsilon$ and the algorithm will stop. Otherwise, $D_{t}\leqslant 4\sqrt{\frac{12}{11}}\|z_{0}-z^{*}\|$. Then, we estimate the additional number of iterations for such bounded $D_{t}$, and finally prove the iteration complexity of Algorithm 2.
@@
Lemma 5
Let $\{z_{k}^{t}\}$ and $\{\bar{z}_{k}^{t}\}$ be the sequences generated by Algorithm 2. At the $t$-th iteration, for all $k\geqslant 1$, we have
@@
Proof  By (38) in Algorithm 2, we get
$(z_{k}^{t})^{*}=$ $\leavevmode\nobreak\ \arg\min_{x}\max_{y}f(x,y)+\dfrac{\sigma_{k}^{t}}{2}\|x-\bar{x}_{k}^{t}\|^{2}-\dfrac{\sigma_{k}^{t}}{2}\|y-\bar{y}_{k}^{t}\|^{2}$
$=$ $\leavevmode\nobreak\ \arg\min_{x}\max_{y}f(x,y)+\dfrac{\sigma_{k-1}^{t}}{2}\|x-\bar{x}_{k-1}^{t}\|^{2}+\dfrac{\sigma_{k}^{t}-\sigma_{k-1}^{t}}{2}\|x-x_{k-1}^{t}\|^{2}$
$\leavevmode\nobreak\ -\dfrac{\sigma_{k-1}^{t}}{2}\|y-\bar{y}_{k-1}^{t}\|^{2}-\dfrac{\sigma_{k}^{t}-\sigma_{k-1}^{t}}{2}\|y-y_{k-1}^{t}\|^{2}$
$=$ $\leavevmode\nobreak\ \arg\min_{x}\max_{y}f_{k-1}^{t}(x,y)+\frac{\gamma_{k}^{t}\sigma_{k}^{t}}{2}\|x-x_{k-1}^{t}\|^{2}-\frac{\gamma_{k}^{t}\sigma_{k}^{t}}{2}\|y-y_{k-1}^{t}\|^{2}.$ (44)
By the optimality condition of $(z_{k}^{t})^{*}$ in (3.1), we then have
@@
Under the convex-concave setting, $f_{k-1}^{t}(x,y)+\frac{\gamma_{k}^{t}\sigma_{k}^{t}}{2}\|x-x_{k-1}^{t}\|^{2}-\frac{\gamma_{k}^{t}\sigma_{k}^{t}}{2}\|y-y_{k-1}^{t}\|^{2}$ is a $\gamma_{k}^{t}\sigma_{k}^{t}$ strongly convex-strongly concave function. Then, by further combining with (45), we have
$\leavevmode\nobreak\ \gamma_{k}^{t}\sigma_{k}^{t}\|(z_{k-1}^{t})^{*}-(z_{k}^{t})^{*}\|^{2}$
$\leqslant$ $\leavevmode\nobreak\ \left\langle F_{k-1}^{t}\big{(}(z_{k-1}^{t})^{*}\big{)}+\gamma_{k}^{t}\sigma_{k}^{t}\big{(}(z_{k-1}^{t})^{*}-z_{k-1}^{t}\big{)}-F_{k-1}^{t}\big{(}(z_{k}^{t})^{*}\big{)}-\gamma_{k}^{t}\sigma_{k}^{t}\big{(}(z_{k}^{t})^{*}-z_{k-1}^{t}\big{)},\right$.
$\leavevmode\nobreak\ \left.(z_{k-1}^{t})^{*}-(z_{k}^{t})^{*}\right\rangle$
$=$ $\leavevmode\nobreak\ \left\langle F_{k-1}^{t}\big{(}(z_{k-1}^{t})^{*}\big{)}+\gamma_{k}^{t}\sigma_{k}^{t}\big{(}(z_{k-1}^{t})^{*}-z_{k-1}^{t}\big{)},(z_{k-1}^{t})^{*}-(z_{k}^{t})^{*}\right\rangle.$ (46)
@@
$\leavevmode\nobreak\ \gamma_{k}^{t}\sigma_{k}^{t}\|(z_{k-1}^{t})^{*}-(z_{k}^{t})^{*}\|^{2}$
$\leqslant$ $\leavevmode\nobreak\ \gamma_{k}^{t}\sigma_{k}^{t}\left\langle(z_{k-1}^{t})^{*}-z_{k-1}^{t},(z_{k-1}^{t})^{*}-(z_{k}^{t})^{*}\right\rangle$
$=$ $\leavevmode\nobreak\ \frac{\gamma_{k}^{t}\sigma_{k}^{t}}{2}\left(\|(z_{k-1}^{t})^{*}-z_{k-1}^{t}\|^{2}+\|(z_{k-1}^{t})^{*}-(z_{k}^{t})^{*}\|^{2}-\|z_{k-1}^{t}-(z_{k}^{t})^{*}\|^{2}\right),$
@@
$\sigma_{k}^{t}=\sum_{i=1}^{k}\alpha_{i}^{t},\leavevmode\nobreak\ \sigma_{k}^{t}\bar{z}_{k}^{t}=\sum_{i=1}^{k}\alpha_{i}^{t}z_{i-1}^{t}.$ (49)
@@
$\leavevmode\nobreak\ \sigma_{k}^{t}\|\bar{z}_{k}^{t}-(z_{k}^{t})^{*}\|$
$\leqslant$ $\leavevmode\nobreak\ \sigma_{k}^{t}\|z_{k-1}^{t}-(z_{k}^{t})^{*}\|+\sigma_{k-1}^{t}\|\bar{z}_{k-1}^{t}-(z_{k-1}^{t})^{*}\|+\sigma_{k-1}^{t}\|(z_{k-1}^{t})^{*}-z_{k-1}^{t}\|$
$\leqslant$ $\leavevmode\nobreak\ (\sigma_{k-1}^{t}+\sigma_{k}^{t})\|z_{k-1}^{t}-(z_{k-1}^{t})^{*}\|+\sigma_{k-1}^{t}\|\bar{z}_{k-1}^{t}-(z_{k-1}^{t})^{*}\|.$
Then, (43) is proved by combining with $\sigma_{0}^{t}=0$.
The following lemma gives an upper bound of the gradient norm at the point $z_{k}^{t}$ by using the distances between $z_{i}^{t}$ and $(z_{i}^{t})^{*}$, $i=0,\cdots,k$.
Lemma 6
Suppose that Assumptions 1 and 3 hold. Let $\{z_{k}^{t}\}$ and $\{\bar{z}_{k}^{t}\}$ be the sequences generated by Algorithm 2. At the $t$-th iteration, for all $k\geqslant 1$, we have
$\|F(z_{k}^{t})\|\leqslant$ $\leavevmode\nobreak\ \bigg{(}\sqrt{\frac{12}{11}}+\frac{72}{11}\bigg{)}M_{k}^{t}\|z_{k}^{t}-(z_{k}^{t})^{*}\|^{2}+2\sqrt{\frac{12}{11}}M_{k}^{t}\bigg{(}\sum_{i=1}^{k-1}\|z_{i}^{t}-(z_{i}^{t})^{*}\|\bigg{)}\|z_{k}^{t}-(z_{k}^{t})^{*}\|$
$\leavevmode\nobreak\ +\sqrt{\frac{12}{11}}M_{k}^{t}\|z_{0}^{t}-(z_{0}^{t})^{*}\|\|z_{k}^{t}-(z_{k}^{t})^{*}\|+\bigg{(}\sqrt{\frac{12}{11}}+1\bigg{)}\sigma_{k}^{t}\|z_{k}^{t}-(z_{k}^{t})^{*}\|$
$\leavevmode\nobreak\ +\sqrt{\frac{12}{11}}\|DF(z_{0}^{t})\|\|z_{k}^{t}-(z_{k}^{t})^{*}\|+\sigma_{1}^{t}\|z_{0}^{t}-(z_{0}^{t})^{*}\|$
$\leavevmode\nobreak\ +\sum_{i=2}^{k}(\sigma_{i-1}^{t}+\sigma_{i}^{t})\|z_{i-1}^{t}-(z_{i-1}^{t})^{*}\|.$ (51)
Proof  By Step 3(d) in Algorithm 2, we have
@@
Under the convex-concave setting, by the optimality condition of $(z_{k}^{t})^{*}$, we further have
$0\leqslant$ $\leavevmode\nobreak\ \langle F_{k}^{t}(\tilde{z}),\tilde{z}-(z_{k}^{t})^{*}\rangle$
$=$ $\leavevmode\nobreak\ \langle F_{k}^{t}(\tilde{z})-F_{k}^{t}(z_{k}^{t})-DF_{k}(z_{k})(\tilde{z}-z_{k}^{t}),\tilde{z}-(z_{k}^{t})^{*}\rangle$
$\leavevmode\nobreak\ +\langle F_{k}^{t}(z_{k}^{t})+DF_{k}^{t}(z_{k}^{t})(\tilde{z}-z_{k}^{t}),\tilde{z}-(z_{k}^{t})^{*}\rangle$
$\leqslant$ $\leavevmode\nobreak\ \frac{M_{k}^{t}}{2}\|\tilde{z}-z_{k}^{t}\|^{2}\|\tilde{z}-(z_{k}^{t})^{*}\|-6M_{k}^{t}\|\tilde{z}-z_{k}^{t}\|\langle\tilde{z}-z_{k}^{t},\tilde{z}-(z_{k}^{t})^{*}\rangle,$
@@
$0\leqslant$ $\leavevmode\nobreak\ \frac{1}{2}\|\tilde{z}-z_{k}^{t}\|^{2}+\frac{1}{2}\|\tilde{z}-(z_{k}^{t})^{*}\|^{2}-6\big{(}\|\tilde{z}-z_{k}^{t}\|^{2}+\|\tilde{z}-(z_{k}^{t})^{*}\|^{2}-\|z_{k}^{t}-(z_{k}^{t})^{*}\|^{2}\big{)}$
$=$ $\leavevmode\nobreak\ -\frac{11}{2}\|\tilde{z}-z_{k}^{t}\|^{2}-\frac{11}{2}\|\tilde{z}-(z_{k}^{t})^{*}\|^{2}+6\|z_{k}^{t}-(z_{k}^{t})^{*}\|^{2},$
@@
$\|F_{k}^{t}(z_{k}^{t})\|=$ $\leavevmode\nobreak\ \|DF_{k}^{t}(z_{k}^{t})(\tilde{z}-z_{k}^{t})+6M_{k}^{t}\|\tilde{z}-z_{k}^{t}\|(\tilde{z}-z_{k}^{t})\|$
$\leqslant$ $\leavevmode\nobreak\ \|DF_{k}^{t}(z_{k}^{t})\|\|\tilde{z}-z_{k}^{t}\|+6M_{k}^{t}\|\tilde{z}-z_{k}^{t}\|^{2}$
$\leqslant$ $\leavevmode\nobreak\ \sqrt{\frac{12}{11}}\|DF_{k}^{t}(z_{k}^{t})\|\|z_{k}^{t}-(z_{k}^{t})^{*}\|+\frac{72M_{k}^{t}}{11}\|z_{k}^{t}-(z_{k}^{t})^{*}\|^{2}.$ (56)
By combining (3.1) with $DF_{k}^{t}(z_{k}^{t})=DF(z_{k}^{t})+\sigma_{k}^{t}I_{m+n}$ and $\frac{\|DF(z_{k}^{t})-DF(z_{0}^{t})\|}{\|z_{k}^{t}-z_{0}^{t}\|}\leqslant M_{k,i}^{t}\leqslant M_{k}^{t}$ in Step 3(d) of Algorithm 2, we have
$\|F_{k}^{t}(z_{k}^{t})\|\leqslant$ $\leavevmode\nobreak\ \sqrt{\frac{12}{11}}\|DF(z_{k}^{t})+\sigma_{k}^{t}I_{m+n}-DF(z_{0}^{t})+DF(z_{0}^{t})\|\|z_{k}^{t}-(z_{k}^{t})^{*}\|$
$\leavevmode\nobreak\ +\frac{72M_{k}^{t}}{11}\|z_{k}^{t}-(z_{k}^{t})^{*}\|^{2}$
$\leqslant$ $\leavevmode\nobreak\ \sqrt{\frac{12}{11}}M_{k}^{t}\|z_{k}^{t}-z_{0}^{t}\|\|z_{k}^{t}-(z_{k}^{t})^{*}\|+\sqrt{\frac{12}{11}}\big{(}\sigma_{k}^{t}+\|DF(z_{0}^{t})\|\big{)}\|z_{k}^{t}-(z_{k}^{t})^{*}\|$
$\leavevmode\nobreak\ +\frac{72M_{k}^{t}}{11}\|z_{k}^{t}-(z_{k}^{t})^{*}\|^{2}.$ (57)
@@
$\|F(z_{k}^{t})\|\leqslant$ $\leavevmode\nobreak\ \sqrt{\frac{12}{11}}M_{k}^{t}\|z_{k}^{t}-z_{0}^{t}\|\|z_{k}^{t}-(z_{k}^{t})^{*}\|+\sqrt{\frac{12}{11}}(\sigma_{k}^{t}+\|DF(z_{0}^{t})\|)\|z_{k}^{t}-(z_{k}^{t})^{*}\|$
$\leavevmode\nobreak\ +\frac{72M_{k}^{t}}{11}\|z_{k}^{t}-(z_{k}^{t})^{*}\|^{2}+\sigma_{k}^{t}\|z_{k}^{t}-(z_{k}^{t})^{*}\|+\sigma_{k}^{t}\|\bar{z}_{k}^{t}-(z_{k}^{t})^{*}\|.$ (58)
@@
$\leavevmode\nobreak\ \|z_{k}^{t}-z_{0}^{t}\|$
$=$ $\leavevmode\nobreak\ \|z_{k}^{t}-(z_{k}^{t})^{*}+(z_{k}^{t})^{*}-z_{k-1}^{t}+z_{k-1}^{t}-(z_{k-1}^{t})^{*}+\cdots+z_{1}^{t}-(z_{1}^{t})^{*}+(z_{1}^{t})^{*}-z_{0}^{t}\|$
$\leqslant$ $\leavevmode\nobreak\ \|z_{k}^{t}-(z_{k}^{t})^{*}\|+\sum_{i=1}^{k-1}2\|z_{i}^{t}-(z_{i}^{t})^{*}\|+\|z_{0}^{t}-(z_{0}^{t})^{*}\|.$ (59)
By plugging (3.1) and (43) into (3.1), the proof is then completed.
@@
Lemma 7
Suppose that Assumptions 1 and 3 hold. Let $\{z_{k}^{t}\}$ and $\{\bar{z}_{k}^{t}\}$ be the sequences generated by Algorithm 2. Set
$\sigma_{k}^{t}=\frac{\epsilon}{41D_{t}}4^{k},N_{k}^{t}=\left\lceil\left(\frac{33\sqrt{3}8^{3-k}H_{k}^{t}D_{t}}{\sigma_{k}^{t}}\right)^{2/3}\right\rceil,$
@@
At the $t$-th iteration, if $D_{t}\geqslant\|z_{0}-z^{*}\|$, then we have $\|F(z_{t+1})\|\leqslant\epsilon$, and $\|z_{t+1}-z^{*}\|\leqslant\frac{135}{56}\|z_{0}-z^{*}\|$, where $z^{*}=(x^{*},y^{*})$ is an optimal solution of (1).
Proof  By Step 3(c), $(x_{k}^{t},y_{k}^{t},H_{k}^{t})$ is an output of the LF-CR algorithm within $N_{k}^{t}$ iterations and $f_{k}^{t}$ is $\rho$-Hessian Lipschitz continuous. We consider two output case of Algorithm 1, i.e., $\|F_{k}^{t}(z_{k}^{t})\|\neq 0$ or $\|F_{k}^{t}(z_{k}^{t})\|=0$. For the case $\|F_{k}^{t}(z_{k}^{t})\|\neq 0$, substituting $f=f_{k}^{t},z=(z_{k}^{t})^{*}$, $z^{*}=(z_{k}^{t})^{*},z_{0}=z_{k-1}^{t},N(\epsilon)-1=N_{k}^{t},H_{N(\epsilon)-1}=H_{k}^{t}$, and $\bar{z}_{N(\epsilon)-1}=z_{k}^{t}$ into (34), we have
$f_{k}^{t}\big{(}x_{k}^{t},(y_{k}^{t})^{*}\big{)}-f_{k}^{t}\big{(}(x_{k}^{t})^{*},y_{k}^{t}\big{)}\leqslant\frac{33\sqrt{3}H_{k}^{t}\|z_{k-1}^{t}-(z_{k}^{t})^{*}\|^{3}}{(N_{k}^{t})^{3/2}}.$ (60)
Under the convex-concave setting, $f_{k}^{t}$ is $\sigma_{k}^{t}$ strongly convex-strongly concave. Then, we further have
$\|z_{k}^{t}-(z_{k}^{t})^{*}\|^{2}$ $\leqslant\frac{33\sqrt{3}H_{k}^{t}\|z_{k-1}^{t}-(z_{k}^{t})^{*}\|^{3}}{\sigma_{k}^{t}(N_{k}^{t})^{3/2}}\leqslant\frac{33\sqrt{3}H_{k}^{t}\|z_{k-1}^{t}-(z_{k-1}^{t})^{*}\|^{3}}{\sigma_{k}^{t}(N_{k}^{t})^{3/2}},$ (61)
where the second inequality is by (42). For the case $\|F_{k}^{t}(z_{k}^{t})\|=0$, we know that $z_{k}^{t}$ = $(z_{k}^{t})^{*}$, so (61) obviously holds.
Next, by (61) and induction, we show that for any $k\geqslant 1$, we have
@@
which implies that $\|z_{k+1}^{t}-(z_{k+1}^{t})^{*}\|\leqslant\frac{1}{8}\|z_{k}^{t}-(z_{k}^{t})^{*}\|$. Next, we further bound the norm $\|F(z_{k}^{t})\|$ in (6) by using (62). By (62) and the choice of $\sigma_{k}^{t}$, we have